OpenStudy (anonymous):

If a + b + c = -9a+b+c=−9 and x + y + z = -3x+y+z=−3, what is 8x + 8z + 8c + 8y + 8a + 8b8x+8z+8c+8y+8a+8b?

OpenStudy (dayakar):

a+b+c= -9--------(I)given x+y+z= -3 -------(II)

OpenStudy (dayakar):

arrange like terms together 8x+8x+8y+8y+8z+8z+8a+8a+8b+8b+8c+8c

OpenStudy (dayakar):

=16x+16y+16z+16a+16b+16c

OpenStudy (dayakar):

=16(x+y+z)+16(a+b+c) taking 16 as common

OpenStudy (dayakar):

=16* (-3) +16 * (-9) according to (I) and (II) =-48-144 =-192

OpenStudy (dayakar):

is it clear